• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Snyder

    Log entry time 23:18:38 on July24,2005

    Entry number 149338

    keyword=VDC reboot

    We asked MCC for 0.4 uA to start HRS momentum scan with the Helium target. The 
    beam current drifted up to 3 uA shortly after this request and as a 
    consequence, the VDC's tripped off.
    
    I paged Rob F., and he directed me to the Crate Resets screen off of the Hall A 
    General Tools. I reset the 2 VDC high voltage lines on both the left and right 
    detector stacks. The VDC's repsonded as we had hoped, and are all now ramping 
    up to 4.0 V